Solve the following ODE’S. Please show your work
f) y +1= (x + y)m - (x + y) + (x + y)P Hint : z = x + y; y = Z - 1
Answer #1

6 7+1 = (x+y) (x+y) + (x+yje - Let us use the transformation =x+y Then dr = 17 de Then from (i), we have 2m 2h+ ZP d2 zm. 1